Your client just retired and is ready to enjoy his hard work. At the moment the. balance in your client's retirement account is $560,000. You expect that you can help the client achieve annual returns of 3.6% in the retirement account during the retirement years. How much money, rounded to the nearest dollar, can your client withdraw from the account MONTHLY in order for the funds in the account to last a total of 25 years?

$2,862

$2,834

$2.965

$2,821
Answer #1

Solution: The answer is $2,834.

Explanation: According to the question

Investment = Present Value = $560,000

Rate of investment = 3.6 %

Rate of interest monthly = 3.6 % / 12 = 0.3 % = 0.003

Number of periods = 25 years

= 25 * 12 = 300 months

Monthly payment = r(PV) / [1-(1+r)-n]

where r = 0.003

PV = $560,000

n = 300 months

Put the values in the formula

Monthly payment = 0.003($560,000) / [1-(1+0.003)-300]

= $1,680 / [1-(1.003)-300]

= $1,680 / [1- 0.40712]

= $1,680 / 0.59288

Monthly payment = $ 2,834 (rounded off)

Please rate positive and comment in case of any doubt. I would be happy to help you further.